A system is defined by the wavefunction ψ(x) = Acos(2πx/L) for –L/4 ≤ x ≤ L/4.

Determine the normalization constant A and the probability that the particle will be

Found between x = 0 and x = L/8?


1
Expert's answer
2021-03-21T11:25:22-0400

The normalization condition is the following (see http://farside.ph.utexas.edu/teaching/qmech/Quantum/node34.html):


"\int_{-L/4}^{L/4}\psi^2(x)dx = 1"

Substituting the wavefunction, obtain:


"\int_{-L/4}^{L/4}A^2\cos^2(2\pi x/L)dx = 1\\"

Taking the integral, find:


"A^2\left( \dfrac{L}{2\pi}\right)\dfrac{\pi }{2} = 1\\
A = \dfrac{2}{\sqrt{L}}"

The probability of finding the particle between 0 and L/8 is:


"\int_{0}^{L/8}\psi^2(x)dx =\int_{0}^{L/8}A^2\cos^2(2\pi x/L)dx = \dfrac{1}{4} + \dfrac{1}{2\pi} \approx 0.409"

Answer. "A = \dfrac{2}{\sqrt{L}},p = 0.409".